I am trying to figure this out, but I don't know the right equation to use. Owners of a candy store want to sell a mix of raisins & nuts for $6 per lb. Raisins usually sell for $3lb., and nuts usually sell for $8lb. They have a 50lb barrel of the raisins. How many lbs of nuts should they mix with raisins so that they hit their target value of $6 per pound for the mixture?
Thanks for your help!

Answer by stanbon(57412)

Owners of a candy store want to sell a mix of raisins & nuts for $6 per lb. Raisins usually sell for $3lb., and nuts usually sell for $8lb. They have a 50lb barrel of the raisins. How many lbs of nuts should they mix with raisins so that they hit their target value of $6 per pound for the mixture?
-------------------------------------------------------
Value Equation:
value of raisins + value of nuts = value of the mixture
3(50) + 8x = 6(50+x)
150 + 9x = 300 + 6x
3x = 150
x = 50 lb (amt of nuts needed for the mixture)
=======================================================
Cheers,
Stan H.